A sinusoidal electromagnetic wave of frequency 40.0 MHz travels in free space in the x direction, determine the wavelength and period of the wave.  


1
Expert's answer
2021-05-08T14:38:09-0400

(a)

"\lambda=\dfrac{c}{f},""\lambda=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{4\cdot10^7\ s^{-1}}=7.5\ m."

(b)

"T=\dfrac{1}{f},""T=\dfrac{1}{4\cdot10^7\ s^{-1}}=25\cdot10^{-9}\ s=25\ ns."